In a population of 1000
Individual 160 individual are affected with thalassemia which is an autosomal recessive disorder find out homozygous and heterozygous dominant individual in the population

Grade:12

1 Answers

Mahima Kanawat
5 years ago
Dear student 
Frequency of recessive individual (p2) is 160/1000 = 0.16,
frequency of recessive allele, p is 0.4
Frequency of dominant allele q is 0.6 and frequency of dominant individuals is q2 =0.36
Therefore  , number of homozygous dominant individuals are 0.36*1000=360
Frequency of heterozygous individual is 2pq = 0.48
Therefore, number of heterozygous dominant individuals are 0.48*1000=480
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT